OpenStudy (anonymous):

x-3 over 3 = x over 2

OpenStudy (ibbutibbu.):

is this the equation you are solving? \[\frac{ x-3 }{ 3 }=\frac{ x }{ ? }\]

OpenStudy (ibbutibbu.):

i mean x/2

OpenStudy (jdoe0001):

\(\bf \cfrac{ x-3 }{ 3 }=\cfrac{ x }{ 2}\quad \textit{try multiplying both sides by }6\)

OpenStudy (anonymous):

or cross multiply and solve for x 2(x-3) = 3x